OpenStudy (anonymous):

solve -2x+1 = 6x = 15

OpenStudy (dumbcow):

get the x's to one side -2x +1 = 6x+15 +2x +2x -------------- 1 = 8x +15 -15 -15 ------------- -14 = 8x Divide by 8 on both sides -14/8 = x Reduce -7/4 = x